OpenStudy (anonymous):

Solve this equation for A: A divided by 2=4

OpenStudy (anonymous):

wats the equation?? or does It mean 2 divided by 4, or 2/4, which is 0.5 or a half (1/2)

OpenStudy (anonymous):

a=8

OpenStudy (anonymous):

8 divided by 2= 4

OpenStudy (anonymous):

Solve for A: A/2 = 4 Multiply both sides by a constant to simplify the equation. Multiply both sides of A/2 = 4 by 2: (2 A)/2 = 2×4 Cancel common terms in the numerator and denominator of (2 A)/2. (2 A)/2 = 2/2×A = A: A = 2×4 Multiply 2 and 4 together. 2×4 = 8: so A = 8